one car is traveling 30 mph faster than the other. The faster car goes 90 miles in the ssame time the slower car travels 60 miles. How fast are the cars traveling?
Answer by jorel1380(2518)

90/x+30=60/x
90x=60x+1800
30x=1800
x=60 mph
x+30= 90 mph..